OpenStudy (anonymous):

John has 15 more baseball cards than Michael. Together they have a total of 55 cards. How many cards does Michael have?

OpenStudy (anonymous):

he has twenty cards

OpenStudy (yamyam70):

please state how did you arrived to the answer, don't just give out answers :)

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

J+15=M J+M=55 J+(J+15)=55 2J= 40 J=20